• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 17:58:19 on November 12, 2007

    Entry number 214076

    This entry is a followup to: 213498

    keyword=BPMA Attenuator Change LHRS

    During the access to fix the AC in the left detector hut, I swapped the attenuator being used for BPMA antenna's 1, 2, and 3 to the original one before the activities on Friday, November 9, 2007. See Halog Entry 213194. I checked the offsets between the new attenuator and the original. The offset was set to zero for the new attenuator and about 0.05 V for the original. Swapping back to the original attenuator should fix the offsets seen in Halog entry # Entry 213381.

    A copy of this log entry has been emailed to: camsonne@jlab.org,saha@jlab.org